# Efficient Latent Variable Causal Discovery: Combining Score Search and Targeted Testing

### v2

@@ 5 Optimizing the Final FCI Orientation Rules
A key feature of FCIT is that the final FCI orientation rules [17] are applied repeatedly throughout the refinement process rather than only once at the end. This ensures maximally informative orientations but introduces performance bottlenecks, as several rules are computationally expensive in their standard form. To make FCIT scalable, we optimize the slowest of these rules. Because the same rules are shared across all FCI-style algorithms, the resulting improvements benefit the entire family, including FCI, GFCI, BOSS-FCI, and GRaSP-FCI.
@@ 6.1 Correctness of BOSS-FCI and GRaSP-FCI
BOSS-FCI and GRaSP-FCI inherit the correctness guarantees of the GFCI algorithm [10]. In its original form, GFCI included an additional Possible-D-SEP search step that ensures oracle correctness. Although this step is often omitted in practice, and no counterexamples have been observed even under a $d$-separation oracle, a formal proof of its redundancy remains open. With that step included, BOSS-FCI and GRaSP-FCI are correct for the same reason as GFCI, differing only in the score-based method used for the initial CPDAG.
6.2 Correctness of LV-Lite
By contrast, LV-Lite lacks formal correctness guarantees. It reports only the PAG equivalence class of the DAG returned by BOSS and therefore cannot remove edges to expose latent confounding or orient bidirected edges. Nevertheless, it always produces a structurally well-formed PAG and, as shown in Section 8, performs well on adjacency and orientation accuracy measures.
To provide intuition for why LV-Lite, while incorrect, nevertheless achieves strong performance on PAG accuracy measures, we note the following considerations. It is unsurprising that LV-Lite performs well on arrowhead precision for common adjacencies (AHPC), since it should in principle produce a subset of the arrowheads that exist over the adjacencies shared with the true PAG. Similarly, adjacency precision (AP) tends to be high: by GFCI theory, LV-Lite should add only a small fraction of “extra” adjacencies, primarily to cover colliders present in the true PAG. Arrowhead recall (AHR) is also expected to be good if the hidden arrowheads represent only a small proportion of those in the true PAG.
Taken together with the high accuracy of BOSS for producing estimates of CPDAGs, these and other such effects explain why LV-Lite generally yields graphs that are very close to the true PAG—especially in large models—despite its lack of formal guarantees. [^6] We did not pursue, but plan to in future work, estimating effects of direct intervention.
@@ 6.5 Soundness of Orientations
Proof sketch. FCIT uses exactly the arrow- and tail-complete rule set R1–R10 of Zhang, [17]. Given correct adjacencies and unshielded colliders, Spirtes and Zhang prove each rule sound; hence all orientations in $G^{\prime}$ are valid.
@@ 7 Well-formedness of PAGs and Markov
One observes that FCIT and LV-Lite always return graphs that are well-formed, in the sense that they satisfy strictly the definition of a PAG. (For a simulation comparison of how well these various algorithms recover PAGs, consider Figure 12 and related figures in the Appendix.) We can determine this independently by checking whether the identified Zhang MAG calculated from estimated graph [17] satisfies the definition of a MAG and whether when converted back to a PAG the orientations of the estimated PAG are recovered precisely. That is, we check the following
@@
These three conditions are individually necessary and jointly sufficient for a singly connected graph over measured variables alone to be a structurally well-formed PAG; we can therefore verify well-formedness by checking these conditions directly. [^7] The Tetrad suite of software contains such a check. By construction, FCIT and LV-Lite always return structurally well-formed PAGs, whereas empirically we find that other FCI-style procedures do not always do so.
@@
  • LV-Lite. LV-Lite trivially returns a well-formed PAG by construction, since it outputs the PAG equivalence class of the BOSS output, which is guaranteed to be a DAG (as BOSS is an algorithm that infers a DAG from a particular permutation of the variables).
@@
We next compare FCIT with FCI, GFCI, BOSS-FCI, GRaSP-FCI, and the heuristic LV-Lite using simulations implemented in Tetrad [12], focusing on accuracy and efficiency in the presence of latent variables.
@@ 8.1.1 Simulation Setup
We simulated data from linear Gaussian structural equation models generated from random forward DAGs with latent variables. [^8] As suggested in an earlier footnote, this is a modeling choice in Tetrad. Scores and tests are also available for the multinomial case, the mixed multinomial/Gaussian case, and the mixed multinomial/post-nonlinear case (see [11]. Also, it should be noted that linear Gaussian scores and tests also work well in the linear non-Gaussian regime [2]). Tests are also available for the general non-parametric case, though not scores currently; these are less scalable. The data were produced using a standard linear Gaussian simulation method (SemSimulation in Tetrad) with the following design:
@@ 8.1.2 Evaluation Metrics
All methods achieve very high adjacency precision. FCIT and LV-Lite consistently lead most orientation-precision metrics (AP/AHP, AHPC, arrow-path precision) and always return valid PAGs. Tail-path precision is highest for BOSS-FCI and GFCI at large $N$, with FCIT close behind. FCIT attains the best bidirected latent-path precision at large samples, while LV-Lite—by design—does not orient bidirected edges. In runtime, LV-Lite (and typically BOSS-FCI) are fastest; GFCI is slowest; FCIT and GRaSP-FCI fall in between. Results for average degrees 2 and 6 (see Appendix) show qualitatively similar patterns, with denser graphs highlighting the tail-path strengths of BOSS-FCI/GFCI and the robustness of FCIT/LV-Lite on PAG validity.
@@
At this larger scale, LV-Lite and FCIT both maintain exceptionally high arrow-path precision ($\approx 0.99$) and consistently produce structurally valid PAGs, though their tail-path precision is somewhat lower ($\approx 0.93$). In contrast, BOSS-FCI and GRaSP-FCI display the opposite trade-off—achieving very high tail-path precision ($\approx 0.97$) but lower arrow-path precision ($\approx 0.89$). Their PAG validity, however, drops to roughly 0.10–0.25. GFCI performs comparably on both arrow and tail paths but with reduced latent-edge precision and moderate PAG validity (0.20), whereas FCI shows the weakest precision scores overall but remains by far the fastest method. Bidirected-edge precision remains low for all methods at this scale, suggesting that in large models such edges should be interpreted cautiously or reserved for qualitative guidance. Overall, these results mirror the smaller-graph findings: FCIT offers the best balance of accuracy, scalability, and structural validity, while LV-Lite remains the most efficient and structurally stable baseline.
@@
Across sparsity levels, FCIT and LV-Lite remain the most accurate and stable overall. Both consistently produce valid PAGs, sustain near-perfect arrow- and tail-path precision, and recover adjacencies reliably even under latent confounding. LV-Lite matches FCIT’s orientation accuracy while running substantially faster, making it a strong baseline when latent structure need not be modeled explicitly. BOSS-FCI and GRaSP-FCI maintain high directional precision but occasionally yield malformed PAGs and incur heavier computational cost at scale. GFCI continues to perform reasonably well on directed edges yet shows moderate PAG validity and weaker latent-edge precision, while standard FCI remains the fastest but the least accurate overall. In sum, FCIT offers the best balance of correctness, scalability, and structural validity among algorithms that consistently return legal PAGs across graph densities.
@@
Figure 1: Adjacency Precision (AP) for 20-node graphs with average degree 4. All algorithms maintain very high adjacency precision across sample sizes. LV-Lite trails slightly. LV-Lite is slightly lower than the others but remains acceptable.
@@
Figure 2: Adjacency Recall (AR) for 20-node graphs with average degree 4. LV-Lite and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover true adjacencies even in the presence of latent confounding. BOSS-FCI, GRaSP-FCI, GFCI, and FCI improve steadily with $N$. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 3: Arrowhead Precision (AHP) for 20-node graphs with average degree 4. LV-Lite and FCIT maintain the highest precision, followed by BOSS-FCI and GRaSP-FCI. FCI remains lowest throughout, while GFCI improves moderately at large NN. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 4: Arrowhead Recall (AHR) for 20-node graphs with average degree 4. FCIT and LV-Lite show the strongest growth in recall with increasing NN, approaching 0.6 at large NN. BOSS-FCI, GFCI, and GRaSP-FCI track closely, while GFCI remains lower overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 5: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 4. LV-Lite and FCIT achieve the highest arrowhead precision, with LV-Lite reaching 0.92 at large NN. BOSS-FCI and GRaSP-FCI show similar trends, while GFCI and especially FCI remain lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 6: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 4. All algorithms improve steadily with increasing NN, with FCI, GRaSP-FCI, and BOSS-FCI converging near 0.9. FCIT levels off slightly below these, while LV-Lite stabilizes around 0.85. GFCI performs comparably to the best methods at large samples. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 7: Tail Precision (TP) for 20-node graphs with average degree 4. GFCI, LV-Lite, and LV-Lite lead here, with other algorihtms lagging behind. This metric measures tail precision relative to the true PAG; for a comparison to the directed paths in the true DAG, see Tail Path Precision.
@@
Figure 8: Tail Recall (TR) for 20-node graphs with average degree 4. LV-Lite and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover correct tail orientations even in the presence of latent confounding. The other algorithms are grouped substantially lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 9: Arrow Path Precision for 20-node graphs with average degree 4. LV-Lite and FCIT again lead in directional precision, followed closely by BOSS-FCI, GRaSP-FCI, and GFCI. FCI trails across sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 12: Proportion of well-formed PAGs for 20-node graphs with average degree 4. As theoretically required, LV-Lite and FCIT achieve perfect PAG validity, whereas other algorithms fall off—sometimes dramatically. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 13: Runtime (CPU milliseconds) for 20-node graphs with average degree 4. All algorithms are quite fast in this regime, with the exception of GFCI, which slows at N=5000N{=}5000. LV-Lite lags behind LV-Lite. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@ 8.2 Real-Data Analysis
We analyzed the Algerian Forest Fire dataset from the UCI Machine Learning Repository [1]. Because we have not discussed nonlinearity or determinism, we removed the six deterministic Canadian Forest Fire System indices—BUI (Build-Up Index), DC (Drought Code), DMC (Duff Moisture Code), FFMC (Fine Fuel Moisture Code), and FWI (Fire Weather Index). We also excluded Day (cyclical variable) and Year (constant at 2012 and therefore non-causal). This left the variables Region (binary), Month, Relative Humidity (RH), Rain, Temperature, Wind Speed (Ws), and Fire (binary occurrence of fire).
Because the dataset mixes discrete and continuous variables, we used the Degenerate Gaussian score and test [2], with penalty discount $1.0$ and $\alpha=0.05$. This method models discrete variables as multinomial and continuous variables as linearly related, enabling hybrid analysis under a unified likelihood framework.
@@
For comparison, LV-Lite returns a PAG nearly identical to the BOSS CPDAG, without introducing bidirected edges. BOSS-FCI and GRaSP-FCI produces similar but slightly denser graphs.
In Ramsey and Andrews, [11], we conducted a complementary analysis using ChatGPT-4 to provide “expert” judgments of edge and orientation correctness for this dataset. This included models where the degenerate Gaussian score was applied without the Canadian Fire System variables (i.e., the same configuration analyzed here), in cases where no domain expert was available. In that study, the FCIT graph was judged to offer the most accurate overall summary, whereas alternative models exhibited spurious adjacencies or misoriented edges. FCIT preserved the accuracy of the BOSS initialization while refining orientations in a manner consistent with PAG semantics, highlighting its advantage—when paired with the degenerate Gaussian test and score—for real-world data containing mixtures of continuous and discrete variables.
Beyond our own benchmark datasets, FCIT and the related nonlinear basis-function test/score [11] have recently been applied to astrophysical data at scale. [6] applied FCIT to a highly nonlinear, non-Gaussian sample of roughly $5\times 10^{5}$ low-redshift galaxies from the NASA Sloan Atlas. Using the Basis-Function LRT for independence testing and Basis-Function BIC for scoring, they recovered a causal graph revealing hierarchical, mass-driven relations among galaxy properties while distinguishing them from observational selection effects, which in this note was judged to be quite plausible, recovering known relationships among the studied variables. This study, submitted to MNRAS, demonstrates that FCIT with an appropriate test and score scales to genuinely astronomical sample sizes and provides interpretable and plausible causal hypotheses for physical mechanisms—marking the first published scientific deployment of FCIT.

### v3

@@ 5 Optimizing the Final FCI Orientation Rules
A key feature of FCIT is that the final FCI orientation rules [zhang2008completeness] are applied repeatedly throughout the refinement process rather than only once at the end. This ensures maximally informative orientations but introduces performance bottlenecks, as several rules are computationally expensive in their standard form. To make FCIT scalable, we optimize the slowest of these rules. Because the same rules are shared across all FCI-style algorithms, the resulting improvements benefit the entire family, including FCI, GFCI, BOSS-FCI, and GRaSP-FCI.
@@ 6.1 Correctness of BOSS-FCI and GRaSP-FCI
BOSS-FCI and GRaSP-FCI inherit the correctness guarantees of the GFCI algorithm [ogarrio2016hybrid]. In its original form, GFCI included an additional Possible-D-SEP search step that ensures oracle correctness. Although this step is often omitted in practice, and no counterexamples have been observed even under a $d$-separation oracle, a formal proof of its redundancy remains open. With that step included, BOSS-FCI and GRaSP-FCI are correct for the same reason as GFCI, differing only in the score-based method used for the initial CPDAG.
6.2 Correctness of LV-Dumb
By contrast, LV-Dumb lacks formal correctness guarantees. It reports only the PAG equivalence class of the DAG returned by BOSS and therefore cannot remove edges to expose latent confounding or orient bidirected edges. Nevertheless, it always produces a structurally well-formed PAG and, as shown in Section 8, performs well on adjacency and orientation accuracy measures.
To provide intuition for why LV-Dumb, while incorrect, nevertheless achieves strong performance on PAG accuracy measures, we note the following considerations. It is unsurprising that LV-Dumb performs well on arrowhead precision for common adjacencies (AHPC), since it should in principle produce a subset of the arrowheads that exist over the adjacencies shared with the true PAG. Similarly, adjacency precision (AP) tends to be high: by GFCI theory, LV-Dumb should add only a small fraction of “extra” adjacencies, primarily to cover colliders present in the true PAG. Arrowhead recall (AHR) is also expected to be good if the hidden arrowheads represent only a small proportion of those in the true PAG.
Taken together with the high accuracy of BOSS for producing estimates of CPDAGs, these and other such effects explain why LV-Dumb generally yields graphs that are very close to the true PAG—especially in large models—despite its lack of formal guarantees. [^6] We did not pursue, but plan to in future work, estimating effects of direct intervention.
@@ 6.5 Soundness of Orientations
Proof sketch. FCIT uses exactly the arrow- and tail-complete rule set R1–R10 of zhang2008completeness. Given correct adjacencies and unshielded colliders, Spirtes and Zhang prove each rule sound; hence all orientations in $G^{\prime}$ are valid.
@@ 7 Well-formedness of PAGs and Markov
One observes that FCIT and LV-Dumb always return graphs that are well-formed, in the sense that they satisfy strictly the definition of a PAG. (For a simulation comparison of how well these various algorithms recover PAGs, consider Figure 12 and related figures in the Appendix.) We can determine this independently by checking whether the identified Zhang MAG calculated from estimated graph [zhang2008completeness] satisfies the definition of a MAG and whether when converted back to a PAG the orientations of the estimated PAG are recovered precisely. That is, we check the following
@@
These three conditions are individually necessary and jointly sufficient for a singly connected graph over measured variables alone to be a structurally well-formed PAG; we can therefore verify well-formedness by checking these conditions directly. [^7] The Tetrad suite of software contains such a check. By construction, FCIT and LV-Dumb always return structurally well-formed PAGs, whereas empirically we find that other FCI-style procedures do not always do so.
@@
  • LV-Dumb. LV-Dumb trivially returns a well-formed PAG by construction, since it outputs the PAG equivalence class of the BOSS output, which is guaranteed to be a DAG (as BOSS is an algorithm that infers a DAG from a particular permutation of the variables).
@@
We next compare FCIT with FCI, GFCI, BOSS-FCI, GRaSP-FCI, and the heuristic LV-Dumb using simulations implemented in Tetrad [ramsey2018tetrad], focusing on accuracy and efficiency in the presence of latent variables.
@@ 8.1.1 Simulation Setup
We simulated data from linear Gaussian structural equation models generated from random forward DAGs with latent variables. [^8] As suggested in an earlier footnote, this is a modeling choice in Tetrad. Scores and tests are also available for the multinomial case, the mixed multinomial/Gaussian case, and the mixed multinomial/post-nonlinear case (see [ramsey2025scalablecausaldiscoveryrecursive]. Also, it should be noted that linear Gaussian scores and tests also work well in the linear non-Gaussian regime [andrews2019learning]). Tests are also available for the general non-parametric case, though not scores currently; these are less scalable. The data were produced using a standard linear Gaussian simulation method (SemSimulation in Tetrad) with the following design:
@@ 8.1.2 Evaluation Metrics
All methods achieve very high adjacency precision. FCIT and LV-Dumb consistently lead most orientation-precision metrics (AP/AHP, AHPC, arrow-path precision) and always return valid PAGs. Tail-path precision is highest for BOSS-FCI and GFCI at large $N$, with FCIT close behind. FCIT attains the best bidirected latent-path precision at large samples, while LV-Dumb—by design—does not orient bidirected edges. In runtime, LV-Dumb (and typically BOSS-FCI) are fastest; GFCI is slowest; FCIT and GRaSP-FCI fall in between. Results for average degrees 2 and 6 (see Appendix) show qualitatively similar patterns, with denser graphs highlighting the tail-path strengths of BOSS-FCI/GFCI and the robustness of FCIT/LV-Dumb on PAG validity.
@@
At this larger scale, LV-Dumb and FCIT both maintain exceptionally high arrow-path precision ($\approx 0.99$) and consistently produce structurally valid PAGs, though their tail-path precision is somewhat lower ($\approx 0.93$). In contrast, BOSS-FCI and GRaSP-FCI display the opposite trade-off—achieving very high tail-path precision ($\approx 0.97$) but lower arrow-path precision ($\approx 0.89$). Their PAG validity, however, drops to roughly 0.10–0.25. GFCI performs comparably on both arrow and tail paths but with reduced latent-edge precision and moderate PAG validity (0.20), whereas FCI shows the weakest precision scores overall but remains by far the fastest method. Bidirected-edge precision remains low for all methods at this scale, suggesting that in large models such edges should be interpreted cautiously or reserved for qualitative guidance. Overall, these results mirror the smaller-graph findings: FCIT offers the best balance of accuracy, scalability, and structural validity, while LV-Dumb remains the most efficient and structurally stable baseline.
@@
Across sparsity levels, FCIT and LV-Dumb remain the most accurate and stable overall. Both consistently produce valid PAGs, sustain near-perfect arrow- and tail-path precision, and recover adjacencies reliably even under latent confounding. LV-Dumb matches FCIT’s orientation accuracy while running substantially faster, making it a strong baseline when latent structure need not be modeled explicitly. BOSS-FCI and GRaSP-FCI maintain high directional precision but occasionally yield malformed PAGs and incur heavier computational cost at scale. GFCI continues to perform reasonably well on directed edges yet shows moderate PAG validity and weaker latent-edge precision, while standard FCI remains the fastest but the least accurate overall. In sum, FCIT offers the best balance of correctness, scalability, and structural validity among algorithms that consistently return legal PAGs across graph densities.
@@
Figure 1: Adjacency Precision (AP) for 20-node graphs with average degree 4. All algorithms maintain very high adjacency precision across sample sizes. LV-Dumb trails slightly. LV-Dumb is slightly lower than the others but remains acceptable.
@@
Figure 2: Adjacency Recall (AR) for 20-node graphs with average degree 4. LV-Dumb and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover true adjacencies even in the presence of latent confounding. BOSS-FCI, GRaSP-FCI, GFCI, and FCI improve steadily with $N$. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 3: Arrowhead Precision (AHP) for 20-node graphs with average degree 4. LV-Dumb and FCIT maintain the highest precision, followed by BOSS-FCI and GRaSP-FCI. FCI remains lowest throughout, while GFCI improves moderately at large NN. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 4: Arrowhead Recall (AHR) for 20-node graphs with average degree 4. FCIT and LV-Dumb show the strongest growth in recall with increasing NN, approaching 0.6 at large NN. BOSS-FCI, GFCI, and GRaSP-FCI track closely, while GFCI remains lower overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 5: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 4. LV-Dumb and FCIT achieve the highest arrowhead precision, with LV-Dumb reaching 0.92 at large NN. BOSS-FCI and GRaSP-FCI show similar trends, while GFCI and especially FCI remain lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 6: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 4. All algorithms improve steadily with increasing NN, with FCI, GRaSP-FCI, and BOSS-FCI converging near 0.9. FCIT levels off slightly below these, while LV-Dumb stabilizes around 0.85. GFCI performs comparably to the best methods at large samples. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 7: Tail Precision (TP) for 20-node graphs with average degree 4. GFCI, LV-Dumb, and LV-Dumb lead here, with other algorihtms lagging behind. This metric measures tail precision relative to the true PAG; for a comparison to the directed paths in the true DAG, see Tail Path Precision.
@@
Figure 8: Tail Recall (TR) for 20-node graphs with average degree 4. LV-Dumb and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover correct tail orientations even in the presence of latent confounding. The other algorithms are grouped substantially lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 9: Arrow Path Precision for 20-node graphs with average degree 4. LV-Dumb and FCIT again lead in directional precision, followed closely by BOSS-FCI, GRaSP-FCI, and GFCI. FCI trails across sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
Figure 12: Proportion of well-formed PAGs for 20-node graphs with average degree 4. As theoretically required, LV-Dumb and FCIT achieve perfect PAG validity, whereas other algorithms fall off—sometimes dramatically. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 13: Runtime (CPU milliseconds) for 20-node graphs with average degree 4. All algorithms are quite fast in this regime, with the exception of GFCI, which slows at N=5000N{=}5000. LV-Dumb lags behind LV-Dumb. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@ 8.2 Real-Data Analysis
We analyzed the Algerian Forest Fire dataset from the UCI Machine Learning Repository [algerian_forest_fires_547]. Because we have not discussed nonlinearity or determinism, we removed the six deterministic Canadian Forest Fire System indices—BUI (Build-Up Index), DC (Drought Code), DMC (Duff Moisture Code), FFMC (Fine Fuel Moisture Code), and FWI (Fire Weather Index). We also excluded Day (cyclical variable) and Year (constant at 2012 and therefore non-causal). This left the variables Region (binary), Month, Relative Humidity (RH), Rain, Temperature, Wind Speed (Ws), and Fire (binary occurrence of fire).
Because the dataset mixes discrete and continuous variables, we used the Degenerate Gaussian score and test [andrews2019learning], with penalty discount $1.0$ and $\alpha=0.05$. This method models discrete variables as multinomial and continuous variables as linearly related, enabling hybrid analysis under a unified likelihood framework.
@@
For comparison, LV-Dumb returns a PAG nearly identical to the BOSS CPDAG, without introducing bidirected edges. BOSS-FCI and GRaSP-FCI produces similar but slightly denser graphs.
In ramsey2025scalablecausaldiscoveryrecursive, we conducted a complementary analysis using ChatGPT-4 to provide “expert” judgments of edge and orientation correctness for this dataset. This included models where the degenerate Gaussian score was applied without the Canadian Fire System variables (i.e., the same configuration analyzed here), in cases where no domain expert was available. In that study, the FCIT graph was judged to offer the most accurate overall summary, whereas alternative models exhibited spurious adjacencies or misoriented edges. FCIT preserved the accuracy of the BOSS initialization while refining orientations in a manner consistent with PAG semantics, highlighting its advantage—when paired with the degenerate Gaussian test and score—for real-world data containing mixtures of continuous and discrete variables.
Beyond our own benchmark datasets, FCIT and the related nonlinear basis-function test/score [ramsey2025scalablecausaldiscoveryrecursive] have recently been applied to astrophysical data at scale. [desmond2025causalstructuregalacticastrophysics] applied FCIT to a highly nonlinear, non-Gaussian sample of roughly $5\times 10^{5}$ low-redshift galaxies from the NASA Sloan Atlas. Using the Basis-Function LRT for independence testing and Basis-Function BIC for scoring, they recovered a causal graph revealing hierarchical, mass-driven relations among galaxy properties while distinguishing them from observational selection effects, which in this note was judged to be quite plausible, recovering known relationships among the studied variables. This study, submitted to MNRAS, demonstrates that FCIT with an appropriate test and score scales to genuinely astronomical sample sizes and provides interpretable and plausible causal hypotheses for physical mechanisms—marking the first published scientific deployment of FCIT.